5.2.4 Suppose a random variable, x, arises from a binomial experiment. If n = 6, and...

5.2.4

Suppose a random variable, x, arises from a binomial experiment. If n = 6, and p = 0.30, find the following probabilities using technology.

A.) P(x=1)

B.) P(x=5)

C.) P(x=3)

D.) P(x \leq 3)

E.) P(x\geq5)

F.) P(x\leq4)
Answer #1

X ~ binomial (n,p)

Where n = 6 , p = 0.30

Binomial probability distribution is

P(X) = nCx px (1-p)n-x

A)

P (X = 1) = 6C1 0.301 0.705

= 0.3025

b)

P (X = 5) = 6C5 0.305 0.701

= 0.0102

c)

P (X = 3) = 6C3 0.303 0.703

= 0.1852

d)

P( X <= 3) = 1 - P (X >= 4)

= 1 - [ P (X = 4) + P (X = 5) + P( X = 6) ]

= 1 - [ 6C4 0.304 0.702 +6C5 0.305 0.70 +6C6 0.306 0.700 ]

= 1 - 0.0705

= 0.9295

e)

P (X >= 5) = P( X = 5) + P (X = 6)

= 6C5 0.305 0.70 +6C6 0.306 0.700

= 0.0109

f)

P (X <= 4) = 1 - P( X >= 5)

= 1 - 0.0109

= 0.9891